The sum of the radii of two circle is 140 cm and the different of their circumference is 88 cm. Find the diameters of the circles

Grade:12th Pass

2 Answers

Ravi
9 years ago
r1+r2=140 ...(1)

2.pi.(r1-r2)=88
r1-r2= (88.7)/(2.22)=28
r1-r2=28 ...(2)

Use equation 1 and 2 to solve for r1 and r2.
Find D1=2r1 and D2=2r2
divya
9 years ago
diameters are 126 and 154 resp.